Operations Report – January 5th

Crew 238 Operations Report 05-01-2022


Name of person filing report: Simon Werner

Non-nominal systems: 1 CO- Sensor, lower deck, replacement in progress
(Atila), spacesuits, radio, see below.

Notes on non-nominal systems: Spacesuit’s issues/ radio battery, see below

Spirit rover used: No

Hours: (taken SOL:2): 149.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: No

Hours: 79.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Curiosity rover used: No

Hours: 158.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Perseverance rover used: no

Hours: unknown

Beginning charge: unknown

Ending charge: unknown

Currently charging: unknown

General notes and comments: none

Summary of Hab operations:


Water (static tank): 449

Water (loft tank): 24 gallons, filled to 55 gallons

Water Meter: post pumping 01544462

Static to Loft Pump used – yes

Static tank pipe heater (on or off): on

Static tank heater (On or off) installed, on

Toilet tank emptied: no

Summary of internet: nominal

Summary of suits and radios:.

Spacesuit Nr.8:

Opened it, checked cable connections with multimeter. One cable was
broken between plug (V+ in) and switch. Removed the cable and soldered
a new one to replace it. Cable shoe was recycled. Battery was low,
5.1V, reinstalled all, currently charging, check tomorrow if charging
was successful.

Spacesuit Nr. 2:

Battery was low 5.6V, checked cables and connections to PCB with
multimeter, no obvious problems. Took the battery out and connected to
Battery charger as suggested by Scott. Check tomorrow again. Unit
currently in the RAM, Battery in HAB.

Spacesuit Nr. 4: Still low, fans slow rapidly down, not checked the
internals yet. Note: the LED of the charger is off. Tested the charger
in different power sockets, no change. Before I go for opening the
unit, shall I crosscheck with a different battery charger or open Nr.
4 first and check connections inside, suggestions?

Spacesuit Nr. 5: After charging overnight nominal again.

Spacesuit Nr 11: After charging overnight nominal again.

All other Spacesuits are nominal and connected to the chargers.

Radio issues: Switched the low radio battery to another radio as
suggested by Atila, the issue remains so the battery is broken –>
replacement by Outpost?

Summary of GreenHab operations:

WATER USE: 200 gallons (guess)

Heater: On

Supplemental light: On

Harvest: Yes, see Green Hab report/ notebook in Green Hab

Summary of Science Dome operations: Aga checked the microscope in the
middle for functionality will be used tomorrow.

Dual split: status unknown, operated by outpost

Summary of RAM operations:

Used for fixing Spacesuit Nr. 8 and 2, Soldering iron was used and
several tools. Cleaned up as possible.

Summary of any observatory issues: none

Summary of health and safety issues: none, see HSO report

Questions, concerns, and requests to Mission Support:

Question: is there any cable for repairing the spacesuits available in
general? Also, heat-shrinking insulation tubes and cable shoes could
not be found. Is there something stored in the Outpost? If so, please
provided some, thanks.

Coffee filter re-supply delivered – very important, thank you.


Copyright © The Mars Society. All rights reserved. | Main Site